Create a group with 8 members or ask any 8 people near by. Inquire a specific question that a possible answer is a number. Record those answers of each member of the group or the people you have surveyed. solve for the sample statistic using the computations that we have discussed.

You may also choose from any of the following topics for survey.

  • The number of Philippine festivals you have attended
  • The number of historical places you have visited
  • The number of filipino heroes you have known


1
Expert's answer
2021-06-03T06:55:35-0400

Let the answers question "The number of historical places you have visited " will be:

"11,8,4,20,6,14,8,15"

Then:


The mean:

"\overline{x}=\frac{\sum x_i}{N}=\frac{11+8+4+20+6+14+8+15}{8}=10.75"


The sample standard deviation:

"\sigma=\sqrt{\frac{\sum (x_i-\overline{x})^2}{N-1}}=\sqrt{\frac{0.25^2+2.75^2+6.75^2+9.25^2+4.75^2+3.25^2+2.75^2+4.25^2}{7}}=28.21"


Put the data in the increasing order:

"4,6,8,8,11,14,15,20"

The median (the midpoint of the dataset):

"m=\frac{8+11}{2}=9.5"

The first quartile (the midpoint of the first half of the dataset):

"Q_1=\frac{6+8}{2}=7"

The third quartile (the midpoint of the second half of the dataset):

"Q_3=\frac{14+15}{2}=14.5"


The mode (the element with the highest frequency):

"M=8"